# Semma: A Semantic Aware Knowledge Graph Foundation Model

### v1

@@ 5 Experiments
Through our experiments, we address the following questions: RQ1: Where does textual semantics matter? RQ2: Does adding textual semantics increase average performance? RQ3: Can textual semantics help generalize to newer, harder settings? RQ4: Do the existing benchmarks suffice?
@@ 5.1 Experimental Setup
Datasets, Baseline, and Setup. We use the same setup as Ultra, pretraining on 3 datasets and testing on 54. They are categorized into three categories: Transductive, Partially Inductive (Inductive $e$), and Fully Inductive (Inductive $e,r$), depending on the overlap between $\mathcal{G}_{\textsc{train}}$ and $\mathcal{G}_{\textsc{test}}$ as discussed in Section 2. We use Ultra as the main baseline for fair comparison. We leave it for future work to extend this to Trix and Motif, where our framework can be adapted to fit in their pipeline. Semma is relatively small and has around 227k parameters. All experiments were run on 2 NVIDIA A100 GPUs. More details in Appendices B and D.
@@
Design Choices. We use gpt-4o-2024-11-20 OpenAI et al. (2024) as the LLM, jina-embeddings-v3 Sturua et al. (2024) as the text embedding model, combined_sum for combining the embeddings and a threshold of $0.8$ to construct ${\mathcal{G}_{R}^{\textsc{text}}}$ based on insights from Putra and Tokunaga (2017) and MLP for fusing ${\mathcal{G}_{R}^{\textsc{str}}}$ and ${\mathcal{G}_{R}^{\textsc{text}}}$. Ablations of these choices are discussed in Section 5.3.
@@
Metafam $+0.3303$ $+0.2935$
WN18RRInductive:v1 $+0.1058$ $+0.0550$
ConceptNet100k $+0.0620$ $+0.1188$
YAGO310 $-0.0879$ $-0.0626$
Hetionet $-0.0556$ $-0.0487$
WikiTopicsMT4:sci $-0.0364$ $-0.0709$
@@ 5.2 Where does textual semantics matter?
Before looking at the average numbers across categories, we first analyse the impact of Semma on the dataset level. From Table 3, we can observe that for datasets like Metafam, the performance increase is almost 2x, while there are some datasets like YAGO310, where there is a considerable drop in MRR and H@10. We provide more insights on this in Appendix E. This hints at how textual semantics can be helpful for some datasets, subject to how rich the relation texts are. To leverage this insight, we introduce the Semma  Hybrid variant, where we switch off the text processing module based on the validation set’s performance.
@@ 5.3 Does adding textual semantics increase average performance?
From Table 1, we can see that Semma clearly outperforms Ultra averaged across all 54 datasets. MRR has an increase of 9 points, and H@10 has an increase of 7 points on average. Note that in the domain of KGFMs, increases of this magnitude are significant, as the tasks are quite demanding Zhang et al. (2025b); Huang et al. (2025). We also perform the Mann-Whitney U test to confirm the statistical significance of the performance gains of Semma. Total Avg MRR has a p-value of 0.0040 (< 0.05) across the runs, and Total Avg Hits@10 has a p-value of 0.0278 (< 0.05). We provide dataset-wise results in Appendix G.
@@
How robust is Semma to design choices? We ablate all the components of Semma to study the impact of the design choices we made. We experiment with other cheaper open-weight LLMs: DeepSeek V3 DeepSeek-AI et al. (2025) and Qwen3 32b Yang et al. (2025), SentenceBERT Reimers and Gurevych (2019) as the embedding model, and other choices described in Section 4.3 and Section 4.2. Results from Table 5 reinforce our final design choices, with regime-wise results in Table 10.
@@ 5.4 Can textual semantics help generalize to newer, harder settings?
Similar to the low-resource setting proposed by Wang et al. (2024), we create a harder evaluation setting where the relation vocabulary of the queries is disjoint from $\mathcal{R}_{\textsc{test}}$, inspired by time-evolving KGs (Cai et al., 2024) where new relations are potentially introduced over time. We discuss this setting in detail in Appendix F. We adapt Ultra and Semma to work in this setting by generating a new ${\mathcal{G}_{R}^{\textsc{str}}}$ and ${\mathcal{G}_{R}^{\textsc{text}}}$ for each query. For queries that share the same head but have different relations, the corresponding ${\mathcal{G}_{R}^{\textsc{str}}}$ will be identical, causing Ultra to produce the same predictions for all such queries, whereas ${\mathcal{G}_{R}^{\textsc{text}}}$ will be able to distinguish them based on the relation identifier. For a working example, refer to Fig. 8. As observed in Table 4, Ultra’s performance drops substantially in the new setting where Semma is 2x better.
@@ 5.5 Do the existing benchmarks suffice?
In some datasets, we identified an overlap where triples from the pretraining data appeared in the test data. We call this common set of triples, the “leaked set”. To assess the impact of this, we conduct a study where we remove all datasets exhibiting any leakage from our evaluation and report these “unleaked” results in Table 2, with more dataset-wise leakage statistics reported in Appendix C. From the results in Table 2, Semma maintains its superior performance over Ultra on this cleaner, more challenging subset of 22 KGs. However, this analysis also highlights a broader issue: the current KGFM benchmarks, while diverse in some structural aspects, often draw from a limited set of popular, large-scale KGs (such as Freebase and WordNet) for both pretraining and testing. This homogeneity can lead to an overestimation of true generalization to genuinely novel domains and ontologies not represented in these common sources.
@@ 6 Conclusion and Future Directions
We introduce Semma, a KGFM that utilizes LLM-derived textual semantics for relations alongside structure. By incorporating a textual relation graph, Semma outperforms purely structural baselines such as Ultra across 54 diverse KGs and shows strong generalization in harder settings with unseen relation vocabularies. This work is a first step toward semantically grounded KGFMs. While we focused on relation-level semantics, future works need to explore extending Semma to entity-level text, richer multilingual inputs, and integration into more expressive models like Trix and Motif. Building on our findings regarding current dataset limitations, the domain also needs new benchmarks with KGs from genuinely diverse and unseen domains to rigorously evaluate true generalization.
@@ Acknowledgments
The authors would like to thank Akshit Sinha for helpful feedback and help with the figures. AA was funded by the CHIPS Joint Undertaking (JU) under grant agreement No. 101140087 (SMARTY), and by the German Federal Ministry of Education and Research (BMBF) under the sub-project with the funding number 16MEE0444. AV was supported by the “UNREAL: Unified Reasoning Layer for Trustworthy ML” project (EP/Y023838/1) selected by the ERC and funded by UKRI EPSRC. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting AA. AA also thanks the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. The authors gratefully acknowledge compute time on HoreKa HPC (NHR@KIT), funded by the BMBF and Baden-Württemberg’s MWK through the NHR program, with additional support from the DFG; and on the Artificial Intelligence Software Academy (AISA) cluster funded by the Ministry of Science, Research and Arts of Baden-Württemberg.
@@
AA conceived the project based on discussions with SS and led the overall design of Semma. AA led the setup and experiments of RQ1 (Section 5.2), RQ2 (Section 5.3) with the help of SK. AA and SK wrote the necessary code and ran experiments for all the RQs. AA and MN discovered the data leakage problem and, with the help of AV, proposed and ran experiments for RQ4 (Section 5.5) with help from SK. Based on an initial idea by BX, SK led the setup and experiments for RQ3 (Section 5.4) with the help of AA. AA wrote the initial draft with help from SK, where all other authors actively contributed to refining it. MN, BX, and AV actively advised on the design of all the experiments. SS and PK provided feedback and advice throughout the project.
@@ Appendix A Prompt
We decompose the prompt into two steps - one to extract the cleaned relation names and the second to obtain the relation descriptions. We define clear rules for each of them and specify the output format. The exact prompt is listed in Fig. 5 and the system instruction in Fig. 3.
@@
We use the OpenRouter [^1] https://openrouter.ai/ API to query the LLMs. In total, the experiments were pretty cost-effective, we spent \$8 on openai/gpt-4o-2024-11-20, \$0.8 on deepseek/deepseek-chat-v3-0324, and \$0.67 on qwen/qwen3-32b-04-28, totalling to less than \$10 for all the 57 datasets cumulatively. We also attach a sample output from openai/gpt-4o-2024-11-20, which can be found at Fig. 4.
@@ Appendix B Datasets
We follow the same evaluation strategy as Ultra and conduct experiments on 57 publicly available KGs spanning a variety of sizes and domains. These datasets are grouped into three generalisation regimes: transductive, inductive with new entities ($e$), and inductive with new entities and relations ($e,r$) at test time. The detailed statistics for transductive datasets is presented in Table 6, inductive ($e$) in Table 7 and inductive ($e,r$) in Table 8. Consistent with Ultra, we perform tail-only evaluation on FB15k237_10, FB15k237_20, and FB15k237_50, where predictions are restricted to the form $(h,r,?)$ during test time.
@@
$||\mathcal{E}|$$||\mathcal{F}|$$||\mathcal{E}|$$|\mathcal{F}|$ Test $||\mathcal{E}|$$|\mathcal{F}|$ Test
@@
$||\mathcal{E}|$$||\mathcal{R}|$$||\mathcal{F}|$$||\mathcal{E}|$$||\mathcal{R}|$$|\mathcal{F}|$ Valid $||\mathcal{E}|$$||\mathcal{R}|$$|\mathcal{F}|$ Test
@@ Appendix C Leakage
We conduct a detailed analysis of both types of leakage and quantify their extent across each dataset. The results are presented in Fig. 6 and Fig. 7, which respectively depict the proportion of datasets affected by indirect (test graph) and direct (query triple) leakage. This analysis highlights the importance of careful dataset curation and motivates the need for evaluating models in settings where such overlaps are explicitly mitigated.
@@ Appendix D Implementation Details
Hyperparameters. For the structure processing module (NBFNet operating on ${\mathcal{G}_{R}^{\textsc{str}}}$) and the final entity-level NBFNet, we largely retain the hyperparameter configurations reported by Ultra (see Table 9 for a summary). Hyperparameters specific to Semma’s novel components were determined as follows: the ${\mathcal{G}_{R}^{\textsc{text}}}$ construction threshold was set to 0.8 based on ablation studies and Putra and Tokunaga (2017); the MLP for fusion ($F$) consists of one hidden layer of size 128 with ReLU activation, chosen based on preliminary experiments on a validation subset of the pre-training data. The choice of LLM (gpt-4o) and text embedding model (Jina-embeddings-v3) was also informed by our ablations (Table 5). We only explore late fusion techniques as we want to reserve the option to switch off the text processing module.
@@
$\text{GNN}_{\textsc{str}}$ layers 6
@@
$\text{GNN}_{\textsc{text}}$ layers 6
@@
$\text{GNN}_{\textsc{ent}}$ layers 6
@@ Appendix E Performance variance in Semma
So, from these preliminary analyses, we hypothesize Semma is especially better for datasets where there is a lot of deviance between the structure and the textual semantics, or if the relation text is rich, or if the KG is sparse.
@@ Appendix F Harder setting
In contrast, Semma effectively addresses this limitation. By incorporating semantic embeddings derived from textual descriptions, Semma distinguishes between novel, previously unseen relations, leveraging the rich semantic signals inherent in relation texts. Consequently, Semma achieves significantly better prediction performance compared to purely structural methods as shown in Table 4.
Consider the example shown in Fig. 8, where Ultra is unable to distinguish between agentcollaborateswithagent and competeswith, resulting in identical and incorrect top-10 predictions for both relations. In contrast, Semma successfully differentiates these two relations, as evident from the clearly ordered top-10 predictions that include the correct ground truth. This clearly illustrates Semma’s ability to leverage textual understanding to overcome structural ambiguities.
@@ Appendix G Full results
We provide the regime-wise averages of the dataset for our ablation studies in Table 10. The complete results per data set reporting MRR and Hits@10 of the zero-shot inference of the pre-trained Ultra and Semma model are presented in Table 11 and Table 12.
@@ Appendix H Discussion on Parallel works
REEF Yu et al. (2025) leverages “relation tokens” from textual descriptions to adaptively generate parameters for GNN components via hypernetworks, aiming for effective pre-training and transfer, while Semma fuses distinct structural and textual relation graph representations for link prediction.
SCR Wang et al. (2025) trains on KGs and reformulates general graph tasks into an inductive KG reasoning format to enable transfer, proposing Semantic Conditional Message Passing. SCR focuses on transferring KG reasoning to general graph tasks by reformatting them, whereas Semma focuses on inductive link prediction within KGs using relation semantics.
KumoRFM Fey et al. (2025) is a pre-trained model for in-context learning on general relational databases, designed for zero-shot predictions across diverse enterprise tasks without specific training, using a Predictive Query Language. Insights like the helpfulness of textual semantics (Section 5.3) can potentially help improve such relational foundation models by leveraging the textual information in the database key and table names.
@@
${\mathcal{G}_{R}^{\textsc{text}}}$ Threshold (0.8) $0.355$ $0.515$ $0.448$ $0.585$ $0.322$ $0.473$ $0.377$ $0.529$
Top-x(20%) $0.341$ $0.509$ $0.453$ $0.591$ $0.31$ $0.462$ $0.371$ $0.525$
Table 10: Ablation Studies. Evaluating the impact of different design choices in Semma. We report MRR and Hits@10 across the three evaluation regimes. The table compares (i) different large language models (LLMs), (ii) language encoders for deriving relation embeddings, (iii) fusion mechanisms (MLP vs. attention), (iv) variations in relation textual input, and (v) different strategies for constructing the textual relation graph (${\mathcal{G}_{R}^{\textsc{text}}}$).

### v2

@@ 5 Experiments
Through our experiments, we address the following questions,
  1. RQ1:
    Where does textual semantics matter?
  2. RQ2:
    Does adding textual semantics increase average performance?
  3. RQ3:
    Can textual semantics help generalize to newer, harder settings?
  4. RQ4:
    Do the existing benchmarks suffice?
@@ 5.1 Experimental Setup
Datasets, Baseline, and Setup. We use the same setup as Ultra, pretraining on 3 datasets and testing on 54. They are categorized into three categories: Transductive, Partially Inductive (Inductive $e$), and Fully Inductive (Inductive $e,r$), depending on the overlap between $\mathcal{G}_{\textsc{train}}$ and $\mathcal{G}_{\textsc{test}}$ as discussed in Section˜2. We use Ultra as the main baseline for fair comparison. We leave it for future work to extend this to Trix and Motif, where our framework can be adapted to fit in their pipeline. Semma is relatively small and has around 227k parameters. All experiments were run on 2 NVIDIA A100 GPUs. More details in Appendices˜B and D.
@@
Design Choices. We use gpt-4o-2024-11-20 OpenAI et al. (2024) as the LLM, jina-embeddings-v3 Sturua et al. (2024) as the text embedding model, combined_sum for combining the embeddings and a threshold of $0.8$ to construct $\mathcal{G}_{R}^{\textsc{text}}$} based on insights from Putra and Tokunaga (2017) and MLP for fusing $\mathcal{G}_{R}^{\textsc{str}}$} and $\mathcal{G}_{R}^{\textsc{text}}$}. Ablations of these choices are discussed in Section˜5.3.
@@
Metafam +0.3303+0.3303 +0.2935+0.2935
WN18RRInductive:v1 +0.1058+0.1058 +0.0550+0.0550
ConceptNet100k +0.0620+0.0620 +0.1188+0.1188
YAGO310 0.0879-0.0879 0.0626-0.0626
Hetionet 0.0556-0.0556 0.0487-0.0487
WikiTopicsMT4:sci 0.0364-0.0364 0.0709-0.0709
@@ 5.2 Where does textual semantics matter?
Before looking at the average numbers across categories, we first analyze the impact of Semma on the dataset level. From Table˜3, we can observe that for datasets like Metafam, the performance increase is almost 2x, while there are some datasets like YAGO310, where there is a considerable drop in MRR and H@10. We provide more insights on this in Appendix˜E. This hints at how textual semantics can be helpful for some datasets, subject to how rich the relation texts are. To leverage this insight, we introduce the SemmaHybrid variant, where we switch off the text processing module based on the validation set’s performance.
@@ 5.3 Does adding textual semantics increase average performance?
From Table˜1, we can see that Semma clearly outperforms Ultra averaged across all 54 datasets, with both MRR and H@10 having considerable improvements. Note that in the domain of KGFMs, increases of this magnitude are significant, as the tasks are quite demanding Zhang et al. (2025b); Huang et al. (2025). We also perform the Mann-Whitney U test to confirm the statistical significance of the performance gains of Semma. Total Avg MRR has a p-value of 0.0040 (< 0.05) across the runs, and Total Avg Hits@10 has a p-value of 0.0278 (< 0.05). We provide dataset-wise results in Appendix˜G.
@@
How robust is Semma to design choices? We ablate all the components of Semma to study the impact of the design choices we made. We experiment with other cheaper open-weight LLMs: DeepSeek V3 DeepSeek-AI et al. (2025) and Qwen3 32b Yang et al. (2025), SentenceBERT Reimers and Gurevych (2019) as the embedding model, and other choices described in Section˜4.3 and Section˜4.2. Results from Table˜5 reinforce our final design choices, with regime-wise results in Table˜10.
@@ 5.4 Can textual semantics help generalize to newer, harder settings?
Similar to the low-resource setting proposed by Wang et al. (2024), we create a harder evaluation setting where the relation vocabulary of the queries is disjoint from $\mathcal{R}_{\textsc{test}}$, inspired by time-evolving KGs (Cai et al., 2024) where new relations are potentially introduced over time. We discuss this setting in detail in Appendix˜F. We adapt Ultra and Semma to work in this setting by generating a new $\mathcal{G}_{R}^{\textsc{str}}$} and $\mathcal{G}_{R}^{\textsc{text}}$} for each query. For queries that share the same head but have different relations, the corresponding $\mathcal{G}_{R}^{\textsc{str}}$} will be identical, causing Ultra to produce the same predictions for all such queries, whereas $\mathcal{G}_{R}^{\textsc{text}}$} will be able to distinguish them based on the relation identifier. For a working example, refer to Figure˜8. As observed in Table˜4, Ultra’s performance drops substantially in the new setting where Semma is 2x better.
@@ 5.5 Do the existing benchmarks suffice?
In some datasets, we identified an overlap where triples from the pretraining data appeared in the test data. We call this common set of triples, the “leaked set”. To assess the impact of this, we conduct a study where we remove all datasets exhibiting any leakage from our evaluation and report these “unleaked” results in Table˜2, with more dataset-wise leakage statistics reported in Appendix˜C. From the results in Table˜2, Semma maintains its superior performance over Ultra on this cleaner, more challenging subset of 22 KGs. However, this analysis also highlights a broader issue: the current KGFM benchmarks, while diverse in some structural aspects, often draw from a limited set of popular, large-scale KGs (such as Freebase and WordNet) for both pretraining and testing. This homogeneity can lead to an overestimation of true generalization to genuinely novel domains and ontologies not represented in these common sources.
@@ 6 Conclusion and Future Directions
We introduce Semma, a KGFM that utilizes LLM-derived textual semantics for relations alongside structure. By incorporating a textual relation graph, Semma outperforms purely structural baselines such as Ultra across 54 diverse KGs and shows strong generalization in harder settings with unseen relation vocabularies. This work is a first step toward semantically grounded KGFMs. While we focused on relation-level semantics, future works need to explore extending Semma to entity-level text, richer multilingual inputs, and integration into more expressive models like Trix and Motif. Building on our findings regarding current dataset limitations, the domain also needs new benchmarks with KGs from genuinely diverse and unseen domains to rigorously evaluate true generalization. Finally, investigating how to extend Semma to yield calibrated predictions (Loconte et al., 2023), and going beyond link prediction and answering a variety of complex queries (Ren and Leskovec, 2020; Galkin et al., 2024b; Gregucci et al., 2025; He et al., 2025) are promising future research avenues.
@@ Acknowledgments
The authors would like to thank Akshit Sinha for helpful feedback and help with the figures. AA was funded by the CHIPS Joint Undertaking (JU) under grant agreement No. 101140087 (SMARTY), and by the German Federal Ministry of Education and Research (BMBF) under the sub-project with the funding number 16MEE0444. AV was supported by the “UNREAL: Unified Reasoning Layer for Trustworthy ML” project (EP/Y023838/1) selected by the ERC and funded by UKRI EPSRC. MN acknowledges BMBF support through the ATLAS project (031L0304A). AA thanks the International Max Planck Research School for Intelligent Systems (IMPRS-IS) and the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. The authors gratefully acknowledge compute time on HoreKa HPC (NHR@KIT), funded by the BMBF and Baden-Württemberg’s MWK through the NHR program, with additional support from the DFG; and on the Artificial Intelligence Software Academy (AISA) cluster funded by the Ministry of Science, Research and Arts of Baden-Württemberg.
@@
AA conceived the project based on discussions with SS and led the overall design of Semma. AA led the setup and experiments of RQ1 (Section˜5.2), RQ2 (Section˜5.3) with the help of SK. AA and SK wrote the necessary code and ran experiments for all the RQs. AA and MN discovered the data leakage problem and, with the help of AV, proposed and ran experiments for RQ4 (Section˜5.5) with help from SK. Based on an initial idea by BX, SK led the setup and experiments for RQ3 (Section˜5.4) with the help of AA. AA wrote the initial draft with help from SK, where all other authors actively contributed to refining it. MN, BX, and AV actively advised on the design of all the experiments. SS and PK provided feedback and advice throughout the project.
@@ Appendix A Prompt
We decompose the prompt into two steps - one to extract the cleaned relation names and the second to obtain the relation descriptions. We define clear rules for each of them and specify the output format. The exact prompt is listed in Figure˜5 and the system instruction in Figure˜3.
@@
We use the OpenRouter [^2] https://openrouter.ai/ API to query the LLMs. In total, the experiments were pretty cost-effective, we spent \$8 on openai/gpt-4o-2024-11-20, \$0.8 on deepseek/deepseek-chat-v3-0324, and \$0.67 on qwen/qwen3-32b-04-28, totalling to less than \$10 for all the 57 datasets cumulatively. We also attach a sample output from openai/gpt-4o-2024-11-20, which can be found at Figure˜4.
@@ Appendix B Datasets
We follow the same evaluation strategy as Ultra and conduct experiments on 57 publicly available KGs spanning a variety of sizes and domains. These datasets are grouped into three generalisation regimes: transductive, inductive with new entities ($e$), and inductive with new entities and relations ($e,r$) at test time. The detailed statistics for transductive datasets is presented in Table˜6, inductive ($e$) in Table˜7 and inductive ($e,r$) in Table˜8. Consistent with Ultra, we perform tail-only evaluation on FB15k237_10, FB15k237_20, and FB15k237_50, where predictions are restricted to the form $(h,r,?)$ during test time.
@@
$\mathcal{E}| ||$$\mathcal{F}|$ $\mathcal{E}| ||$$\mathcal{F}|$ Test $\mathcal{E}| ||$$\mathcal{F}|$ Test
@@
$\mathcal{E}| ||$$\mathcal{R}|$ $\mathcal{F}| ||$$\mathcal{E}|$ $\mathcal{R}| ||$$\mathcal{F}|$ Valid $\mathcal{E}| ||$$\mathcal{R}|$ $|\mathcal{F}|$ Test
@@ Appendix C Leakage
We conduct a detailed analysis of both types of leakage and quantify their extent across each dataset. The results are presented in Figure˜6 and Figure˜7, which respectively depict the proportion of datasets affected by indirect (test graph) and direct (query triple) leakage. This analysis highlights the importance of careful dataset curation and motivates the need for evaluating models in settings where such overlaps are explicitly mitigated.
@@ Appendix D Implementation Details
Hyperparameters. For the structure processing module (NBFNet operating on $\mathcal{G}_{R}^{\textsc{str}}$}) and the final entity-level NBFNet, we largely retain the hyperparameter configurations reported by Ultra (see Table˜9 for a summary). Hyperparameters specific to Semma’s novel components were determined as follows: the $\mathcal{G}_{R}^{\textsc{text}}$} construction threshold was set to 0.8 based on ablation studies and Putra and Tokunaga (2017); the MLP for fusion ($F$) consists of one hidden layer of size 128 with ReLU activation, chosen based on preliminary experiments on a validation subset of the pre-training data. The choice of LLM (gpt-4o) and text embedding model (Jina-embeddings-v3) was also informed by our ablations (Table˜5). We only explore late fusion techniques as we want to reserve the option to switch off the text processing module.
@@
$\text{GNN}_{\textsc{str}}$ # layers 6
@@
$\text{GNN}_{\textsc{text}}$ # layers 6
@@
$\text{GNN}_{\textsc{ent}}$ # layers 6
@@ Appendix E Performance variance in Semma
So, from these preliminary analyses, we hypothesize Semma is especially better for datasets where there is a significant deviance between the structure and the textual semantics, or if the relation text is rich, or if the KG is sparse.
Sensitivity Analysis. We observed that the model’s performance is sensitive to the cosine similarity threshold used for edge creation in $\mathcal{G}_{R}^{\textsc{text}}$}. Consequently, we performed a grid search over the values {0.6, 0.7, 0.79, 0.8} as part of our hyperparameter tuning process. Based on this search and prior works Putra and Tokunaga (2017), we selected a final threshold of 0.8 and further validated this choice by conducting a qualitative analysis of the generated edges on a subset of the datasets. Investigating dynamic thresholding techniques for different datasets remains a promising direction for future work.
@@ Appendix F Harder setting
In contrast, Semma effectively addresses this limitation. By incorporating semantic embeddings derived from textual descriptions, Semma distinguishes between novel, previously unseen relations, leveraging the rich semantic signals inherent in relation texts. Consequently, Semma achieves significantly better prediction performance compared to purely structural methods as shown in Table˜4.
Consider the example shown in Figure˜8, where Ultra is unable to distinguish between agentcollaborateswithagent and competeswith, resulting in identical and incorrect top-10 predictions for both relations. In contrast, Semma successfully differentiates these two relations, as evident from the clearly ordered top-10 predictions that include the correct ground truth. This clearly illustrates Semma’s ability to leverage textual understanding to overcome structural ambiguities.
@@ Appendix G Full results
We provide the regime-wise averages of the dataset for our ablation studies in Table˜10. The complete results per data set reporting MRR and Hits@10 of the zero-shot inference of the pre-trained Ultra and Semma model are presented in Table˜11 and Table˜12.
@@ Appendix H Discussion on Parallel works
REEF Yu et al. (2025) leverages “relation tokens” from textual descriptions to adaptively generate parameters for GNN components via hypernetworks, aiming for effective pre-training and transfer, while Semma fuses distinct structural and textual relation graph representations for link prediction. SCR Wang et al. (2025) trains on KGs and reformulates general graph tasks into an inductive KG reasoning format to enable transfer, proposing Semantic Conditional Message Passing. SCR focuses on transferring KG reasoning to general graph tasks by reformatting them, whereas Semma focuses on inductive link prediction within KGs using relation semantics.
KumoRFM Fey et al. (2025) is a pre-trained model for in-context learning on general relational databases, designed for zero-shot predictions across diverse enterprise tasks without specific training, using a Predictive Query Language. Insights like the helpfulness of textual semantics (Section˜5.3) can potentially help improve such relational foundation models by leveraging the textual information in the database key and table names.
@@
$\mathcal{G}_{R}^{\textsc{text}}$}
Threshold (0.8) $0.355$ $0.515$ $0.448$ $0.585$ $0.322$ $0.473$ $0.377$ $0.529$
Top-x% (20%) $0.341$ $0.509$ $0.453$ $0.591$ $0.31$ $0.462$ $0.371$ $0.525$
Table 10: Ablation Studies. Evaluating the impact of different design choices in Semma. We report MRR and Hits@10 across the three evaluation regimes. The table compares (i) different LLMs, (ii) language encoders for deriving relation embeddings, (iii) fusion mechanisms (MLP vs. Attention), (iv) variations in relation textual input, and (v) different strategies for constructing the textual relation graph ($\mathcal{G}_{R}^{\textsc{text}}$}).